OpenStudy (anonymous):

given the piecewise function shown below, select all of the statements that are true

OpenStudy (anonymous):

I don't see a piecewise function @jm6511

OpenStudy (anonymous):

f(x)= {-x+1,x<0} -2, x=0 x^2-1,x>0 select all that are true a. f(4)=7 b. f(1)= 0 c. f(-1)=2 d. f(-2)=0 sorry clicked enter before i could finish :/

OpenStudy (unklerhaukus):

\[f(x)= \begin{cases}-x+1,&x<0\\ -2,& x=0\\ x^2-1,&x>0\end{cases}\qquad\begin{array}{c}\text{case i}\\\text{case ii}\\\text{case iii}\end{array}\]

OpenStudy (unklerhaukus):

can you tell me which case \(f(4)\) is in; i,ii, or iii?